OpenStudy (anonymous):

A coffee mixture has beans that sell for $1.35 per pound and beans that sell for $1.05 per pound. If 120 pounds of beans create a mixture worth $1.15 per pound, how many pounds of each bean is used? 50 lb of $1.35/lb; 70 lb of $1.05/lb 40 lb of $1.35/lb; 80 lb of $1.05/lb 45 lb of $1.35/lb; 75 lb of $1.05/lb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what do I do?

OpenStudy (shane_b):

Not sure I did that right...I'll check it in a sec. Wife calling :P

OpenStudy (anonymous):

lol ok

OpenStudy (shane_b):

Go ahead Hero...I was setting up two simultaneous equations for it but I'm not sure that's correct.

OpenStudy (anonymous):

hero?

OpenStudy (shane_b):

Well, you can do it also by just trying the answers. Multiply them out to get a total and divide that by 120 to see if it = $1.15

OpenStudy (anonymous):

k

hero (hero):

Sorry, my desktop screen just outrageously crashed on me for no apparent reason!

OpenStudy (shane_b):

Look at the second answer:\[ (40*1.35)+(80*1.05)=138\]\[\frac{138}{120}=1.05\]

OpenStudy (shane_b):

Oops..that should have been 1.15 at the end...anyway :)

OpenStudy (anonymous):

Thanks. @hero its okay

OpenStudy (shane_b):

I see the mistake. I think I had it right after all: 1.35x +1.05y=(120)(1.15) x+y=120 I neglected to multiple $1.15 times the total pounds.

OpenStudy (shane_b):

If you solve for x and y, you will get 40 and 80.

OpenStudy (anonymous):

Great! Thanks

OpenStudy (shane_b):

Me and my silly mistakes :) You're welcome...if you need help solving the equations above just so you know how to do it let me know.

OpenStudy (anonymous):

Oh I understand thanks tho

hero (hero):

I was going to post that equation as a hint before my desktop screen crashed for no reason. I spent the entire hour troubleshooting. I think my surge protector died on me. Had to switch to another one.

OpenStudy (anonymous):

oh ok

hero (hero):

Good thing someone was able to help you

OpenStudy (anonymous):

Yup